OpenStudy (anonymous):

will give medal to best answer! The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. C = 38°, a = 19, c = 10

hero (hero):

Did you try simply using Law of Sines to figure it out?

OpenStudy (anonymous):

yes but i didnt understand

hero (hero):

What did you come up with?

OpenStudy (anonymous):

i got sin A = 67 but that isnt in my answer choices No triangle is formed. A = 58.6°, B = 83.4°, b ≈ 15.6 A = 83.4°, B = 58.6°, b ≈ 15.6 The triangle cannot be solved with the Law of Sines.

hero (hero):

I don't understand how you got that. Please show me the steps you took to get A = 67

OpenStudy (anonymous):

i used c/ sin c = a/ sin a 10/sin 37 = 19/sin a and then i divided 19 sin 38 over sin 10 and i got 67.36

hero (hero):

where did you get the 37 from?

OpenStudy (anonymous):

I meant 38, i used 38 in my claculations and still got 67

OpenStudy (anonymous):

calculations

hero (hero):

Are you in degree mode or radian mode?

OpenStudy (anonymous):

i dont know i just typed it in and got 67, cant you check it?

hero (hero):

You must make sure you're in degree mode. That's very important.

hero (hero):

But either way it goes, you're not doing it correctly.

OpenStudy (anonymous):

oh really, i hadnt noticed that im doing it wrong. I need help thats why im on here obviously

hero (hero):

\[\sin A = \frac{19\sin(38^{\circ})}{10} \\\sin A = \]

hero (hero):

Compute the right hand side again and let me know what you get. Make sure you're in degree mode.

OpenStudy (anonymous):

1.169756803

hero (hero):

Okay now \[\sin A = 1.169756803\] So how do we find A?

OpenStudy (anonymous):

a is 19, sin A is 1.16, what do you mean find A?

hero (hero):

A is an angle. Find it

hero (hero):

A is an Angle like theta

hero (hero):

I'm not talking about a = 19. That's a side length

hero (hero):

\[\sin (A^{\circ}) = 1.169756803\] I'm talking about Angle A. Find that.

OpenStudy (anonymous):

how?

hero (hero):

You take the inverse sine of both sides to get: \[\sin^{-1}(A^{\circ}) = \sin^{-1}({1.169756803}) \\A^{\circ} = \sin^{-1}({1.169756803}) \]

hero (hero):

So compute \[\sin^{-1}({1.169756803})\]

hero (hero):

Let me know what you get.

OpenStudy (anonymous):

it says error

hero (hero):

Exactly.

hero (hero):

That's what it should say because no such angle exists.

hero (hero):

And if the angle doesn't exist, then the triangle can't possibly exist.

hero (hero):

\[-1< \sin(\theta) < 1\] means that the value of sine can only be between - or + one. So there's no such thing as \(\sin(\theta) = 1.17\)